A ship heading to a dock at a constant velocity emits a 500.0 Hz sound from its 3.00 W
horn. The ship is also receding from an island which hears a 461 Hz sound from the ship.
The ship is initially 87.5 km from the dock. At this initial position it takes 4.13 minutes for
sound to travel from the ship to the dock.

A)The island has an identical horn. With both horns sounding, a periodic change in volume
is heard on the island, from very quiet to very loud. How many quiet moments occur in a
second?

B)What is the air temperature?

C). How far must the ship travel from its initial position before the intensity of the ship’s
horn is 100 I0 to someone standing on the dock?

D)What will be the frequency of sound heard at the dock when it is first heard?

E)What will be the frequency of sound heard at the dock when the ship is 10.0 m away?
1
Expert's answer
2018-09-24T12:33:08-0400
A)
f=f_2-f_1=500-461=39 Hz.
We have 39 quiet moments in a second.
B) Speed of sound
v=87500/(4.13(60))=353.1 m/s.
Thus, the air temperature:
T=273((353.1/331)^2-1)=37.7℃
C)
I/I_0 =100=(x_0/x)^2
x=x_0/10=8.75 km.
D)
f=f_2=500 Hz.
E)
f=f_1=461 Hz.